The Ace Job Shop is made up of a production facility where machinists work. In a...

  1. The Ace Job Shop is made up of a production facility where machinists work. In a nearby area, the tool crib, these machinists pick up their tools from an attendant at the tool crib. The Ace Job Shop employs both the machinists and the tool crib attendants. On average 8 machinists arrive at the tool crib every hour. Each machinist is paid $18 per hour. The supervisor can staff the crib with a junior attendant who is paid $10.00/hour and can process 10 arrivals per hour OR with a senior attendant who can process 12 arrivals per hour but is paid $16.00/hour. If the supervisor wishes to minimize total operating costs, which attendant should be selected, and what would be the total, estimated hourly cost at the tool crib?
Answer #1

Arrival rate = 8

With junior attendant:

Average number of machinist in the system = Arrival rate/(Service rate-arrival rate) = 8/(10-8) = 4

Total cost = Cost of machinist+cost of junior attendant

= 4*18+10 = $82/hour

With Senior attendant:

Average number of machinist in the system = Arrival rate/(Service rate-arrival rate) = 8/(12-8) = 2

Total cost = Cost of machinist+cost of junior attendant

= 2*18+16 = $52/hour

So to minimize total operating costs, Senior attendant should be selected, and the total, estimated hourly cost at the tool crib=$52/hour